# DoctorAgent-RL: A Multi-Agent Collaborative Reinforcement Learning System for Multi-Turn Clinical Dialogue

### v1

@@ 4.1 Patient Agent Behavior Analysis on MTMedDialog
4.2 Comparative Performance Evaluation of Multi-Model Approaches on MTMedDialog
Experimental results in Table  2 show DoctorAgent-RL achieving a comprehensive average score of 53.9%, demonstrating significant improvements over frontier models, open-source base models, and domain-specific models. More detailed experimental results can be found in Appendix F. The system particularly maintains stable advantages in disease types requiring in-depth consultation, proving its superior ability to simulate doctors’ clinical consultation processes.
Analysis of model dialogue processes reveals key limitations in existing methods’ interaction quality. While all models receive explicit instructions to "ask only one question at a time," frontier models show strict compliance but lack professional medical consultation knowledge, often missing critical symptom information in their questions. Some open-source models (e.g., GLM-4, Mistral) achieve decent comprehensive average scores but improperly combine multiple questions, reflecting inadequate instruction-following capability that affects complex condition diagnosis. The domain-specific model BioMistral exhibits the highest interaction frequency but frequently repeats similar questions due to ineffective planning, resulting in the lowest diagnostic accuracy. These findings demonstrate that question quality - not quantity - determines diagnostic effectiveness.
@@ 4.2 Comparative Performance Evaluation of Multi-Model Approaches on MTMedDialog
4.3 Adaptive Fine-Tuning Strategies for Task-Specific Optimization
@@
4.4 Impact Analysis of Budget of Turns on Diagnosis and Recommendation Performance
@@
5 Conclusion and future work
We propose DoctorAgent-RL, a multi-agent collaborative reinforcement learning framework that establishes an innovative paradigm enabling LLMs to progressively refine diagnoses through proactive questioning. The framework effectively addresses the unrealistic requirement of conventional systems for patients to fully describe symptoms in a single interaction by synergistically integrating a high-fidelity patient agent simulating pathologically consistent symptom expressions with a doctor agent optimizing adaptive questioning strategies through reinforcement learning, guided by an evaluator providing multidimensional clinical reward signals. Experimental results demonstrate that DoctorAgent-RL significantly outperforms traditional methods in diagnostic accuracy. These findings not only mark a paradigm shift in medical AI from static Q&A to dynamic reasoning, but also provide clinicians with interpretable decision support through its proactive questioning strategies enabled by multi-turn interactions, while simultaneously pioneering new technical pathways for patient-led symptom screening.
Future research will focus on advancing multimodal medical reasoning capabilities by integrating heterogeneous data sources such as medical imaging, pathological slides, and real-time physiological signals from wearable devices to enhance the model’s comprehensive understanding of complex conditions. Concurrently, systematic efforts are needed to address data bias and decision transparency, establishing a fairness evaluation framework that accounts for demographic characteristics and disease spectrum variations to ensure model robustness and interpretability across diverse patient populations. Throughout technological iterations, an ethical governance framework for medical AI must be developed through interdisciplinary research, addressing critical issues such as diagnostic accountability, privacy and data security, and doctor-patient trust mechanisms, ultimately achieving a dynamic balance between technological innovation and adherence to medical ethics and patient safety standards.
Appendix
Appendix A Comparison of Diagnostic Consultation Paradigms
This section introduces the comparison of different diagnostic consultation paradigms, as shown in Figure 4.
Figure 4: Comparison of diagnostic consultation paradigms. (A) Single-Turn Consultation: Direct diagnosis based on initial query without follow-up interactions, prone to misdiagnosis due to insufficient symptom clarification. (B) SFT Multi-Turn Consultation: Implements basic multi-turn dialogue through supervised fine-tuning but lacks adaptive optimization. (C) SFT-RL Multi-Turn Consultation: Integrates RL with real-world feedback to dynamically optimize questioning strategies, enabling context-aware symptom investigation (e.g., pain nature, duration, triggers) and reducing diagnostic errors through iterative interaction.
Appendix B Details of MTMedDialog
@@
Appendix C Prompts for Evaluation Metrics
@@
Prompt for Evaluating the Doctor Agent’s Diagnosis and Recommendation Accuracy Task: As a medical expert, evaluate the semantic similarity between the model-generated medical text and the ground truth reference. Score on a 0–5 point scale based on meaning alignment (wording differences are acceptable if meaning matches). Criteria: 5: Identical meaning (different wording okay). 4: Minor wording/detail differences; overall meaning aligned. 3: Partial meaning overlap; important differences exist but core intent is partially shared. 2: Limited meaning overlap; key details or context differ significantly. 1: Minimal meaning overlap; mostly unrelated or only superficially related. 0: Unrelated or completely different meaning (no meaningful semantic connection). Now evaluate: Candidate: "{candidate}" Reference: "{reference}" OUTPUT FORMAT: <think> [Your Thinking Process] </think><answer> [Your Score] </answer>
@@ 1. Information Control Rate (0-1):
"information_control_rate": float, "response_completeness_rate": float, "factual_conflict_rate": float }
@@
Appendix D Prompt for Doctor Agent
@@
Question: (your question). </answer> | <answer>If you believe you have obtained enough information, please only provide diagnosis and recommendations, in this format: Diagnosis: (the patient’s most likely disease or symptoms) Recommendation: (corresponding treatment plan or advice) </answer>
@@
Appendix E Prompt for Patient Agent
@@
If the question is unrelated to your chief complaint, state: "Sorry, I cannot answer this question." If the question is repetitive, reply: "Sorry, you’ve already asked this question." Your chief complaint:
@@
Appendix F Detailed Comparative Performance Evaluation of Multi-Model Approaches on MTMedDialog
@@
Appendix G Detailed Description of the Dialogue Flow in DoctorAgent-RL
This section provides a detailed explanation of the dialogue flow in DoctorAgent-RL, as illustrated in Figure  10. The framework aims to simulate a realistic diagnostic process by coordinating a structured dialogue flow between two agents: the doctor and the patient. This appendix comprehensively describes the mechanisms for dialogue initiation, progression, and termination.
@@ Appendix G Detailed Description of the Dialogue Flow in DoctorAgent-RL
Appendix H Settings of RL training in DoctorAgent-RL
Hardware & Software. All training and evaluation were performed on a system featuring eight NVIDIA A100 80GB PCIe GPUs, an Intel Xeon Platinum 8369B 32-Core Processor, and 1.0 TB of RAM. For supervised fine-tuning, we utilized the LLaMA-Factory framework [41] to fine-tune with LoRA [42]. Our DoctorAgent-RL, is built upon the VERL framework (v0.2) [43] for reinforcement learning with language models, with RAGEN [44] providing the multi-turn RL architecture. We employed vLLM (v0.8.5) [45] for efficient LLM inference and evaluation, PyTorch (v2.4.0) with CUDA 12.4 for deep learning, and Ray for distributed training and serving. Flash Attention 2 [46] was integrated to optimize attention computation.
Hyperparameter. For completeness and reproducibility, all hyperparameters employed in DoctorAgent-RL are detailed in Table 7. We observed that the decoupled clip approach from DAPO [47] significantly enhances exploration during RL training, and thus, we adopted it for our final training process.
Supervised Fine-Tuning learning rate 1e-4
batch size 64
epochs 3
lora_rank 8
DoctorAgent-RL actor learning rate 1e-6
state masking true
kl loss coef 0.001
kl penalty low_var_kl
entropy coeff 0.001
clip high 0.28
clip low 0.2
batch size 128
epochs 1
rollout group size 8
rollout temperature 0.7

### v2

@@ Introduction
Comparative Performance Evaluation of Multi-Model Approaches on MTMedDialog
Experimental results in Table  2 show DoctorAgent-RL achieving a comprehensive average score of 53.9%, demonstrating significant improvements over frontier models, open-source base models, and domain-specific models. More detailed experimental results can be found in Supplementary Material. The system particularly maintains stable advantages in disease types requiring in-depth consultation, proving its superior ability to simulate doctors’ clinical consultation processes.
Analysis of model dialogue processes reveals key limitations in existing methods’ interaction quality. While all models receive explicit instructions to ”ask only one question at a time,” frontier models show strict compliance but lack professional medical consultation knowledge, often missing critical symptom information in their questions. Some open-source models (e.g., GLM-4, Mistral) achieve decent comprehensive average scores but improperly combine multiple questions, reflecting inadequate instruction-following capability that affects complex condition diagnosis. The domain-specific model BioMistral exhibits the highest interaction frequency but frequently repeats similar questions due to ineffective planning, resulting in the lowest diagnostic accuracy. These findings demonstrate that question quality - not quantity - determines diagnostic effectiveness.
@@
Adaptive Fine-Tuning Strategies for Task-Specific Optimization
@@
Impact Analysis of Budget of Turns on Diagnosis and Recommendation Performance
@@
Conclusion
The framework effectively addresses the unrealistic requirement of conventional systems for patients to fully describe symptoms in a single interaction by synergistically integrating a high-fidelity patient agent simulating pathologically consistent symptom expressions with a doctor agent optimizing adaptive questioning strategies through reinforcement learning, guided by an evaluator providing multidimensional clinical reward signals. Experimental results demonstrate that DoctorAgent-RL significantly outperforms traditional methods in diagnostic accuracy. These findings not only mark a paradigm shift in medical AI from static Q&A to dynamic reasoning, but also provide doctors with interpretable decision support through its proactive questioning strategies enabled by multi-turn interactions, while simultaneously pioneering new technical pathways for patient-led symptom screening. More critically, DoctorAgent-RL offers a practical technical approach to addressing the global physician shortage crisis. In resource-constrained primary care settings and emergency triage scenarios, DoctorAgent-RL efficiently handles initial screenings through dynamic multi-turn dialogue, reducing misdiagnosis risks caused by time pressure on doctors, while freeing clinicians from repetitive inquiries to focus on complex critical cases. This pioneers an AI-driven strategy for optimizing medical resource allocation and alleviating structural workforce shortages.
Supplementary Materials
Details of MTMedDialog
@@ Conclusion
Prompts for Evaluation Metrics
@@
Prompt for Evaluating the Doctor Agent’s Diagnosis and Recommendation Accuracy Task: As a medical expert, evaluate the semantic similarity between the model-generated medical text and the ground truth reference. Score on a 0–5 point scale based on meaning alignment (wording differences are acceptable if meaning matches). Criteria: 5: Identical meaning (different wording okay). 4: Minor wording/detail differences; overall meaning aligned. 3: Partial meaning overlap; important differences exist but core intent is partially shared. 2: Limited meaning overlap; key details or context differ significantly. 1: Minimal meaning overlap; mostly unrelated or only superficially related. 0: Unrelated or completely different meaning (no meaningful semantic connection). Now evaluate: Candidate: ”{candidate}” Reference: ”{reference}” OUTPUT FORMAT: <think> [Your Thinking Process] </think><answer> [Your Score] </answer>
@@
Supervised Fine-Tuning learning rate 1e-4
batch size 64
epochs 3
lora_rank 8
DoctorAgent-RL actor learning rate 1e-6
state masking true
kl loss coef 0.001
kl penalty low_var_kl
entropy coeff 0.001
clip high 0.28
clip low 0.2
batch size 128
epochs 1
rollout group size 8
rollout temperature 0.7
@@ 1. Information Control Rate (0-1):
”information_control_rate”: float, ”response_completeness_rate”: float, ”factual_conflict_rate”: float }
@@
Prompt for Doctor Agent
@@
Question: (your question). </answer><answer>If you believe you have obtained enough information, please only provide diagnosis and recommendations, in this format: Diagnosis: (the patient’s most likely disease or symptoms) Recommendation: (corresponding treatment plan or advice) </answer>
@@
Prompt for Patient Agent
@@
If the question is unrelated to your chief complaint, state: ”Sorry, I cannot answer this question.” If the question is repetitive, reply: ”Sorry, you’ve already asked this question.” Your chief complaint:
@@
Detailed Comparative Performance Evaluation of Multi-Model Approaches on MTMedDialog
@@
Detailed Description of the Dialogue Flow in DoctorAgent-RL
This section provides a detailed explanation of the dialogue flow in DoctorAgent-RL, as illustrated in Figure  9. The framework aims to simulate a realistic diagnostic process by coordinating a structured dialogue flow between two agents: the doctor and the patient. This appendix comprehensively describes the mechanisms for dialogue initiation, progression, and termination.
@@
Settings of RL training in DoctorAgent-RL
Hardware & Software. All training and evaluation were performed on a system featuring eight NVIDIA A100 80GB PCIe GPUs, an Intel Xeon Platinum 8369B 32-Core Processor, and 1.0 TB of RAM. For supervised fine-tuning, we utilized the LLaMA-Factory framework (Zheng et al. 2024) to fine-tune with LoRA (Hu et al. 2022). Our DoctorAgent-RL, is built upon the VERL framework (v0.2) (Sheng et al. 2024) for reinforcement learning with language models, with RAGEN (Wang et al. 2025) providing the multi-turn RL architecture. We employed vLLM (v0.8.5) (Kwon et al. 2023) for efficient LLM inference and evaluation, PyTorch (v2.4.0) with CUDA 12.4 for deep learning, and Ray for distributed training and serving. Flash Attention 2 (Dao 2023) was integrated to optimize attention computation.
Hyperparameter. For completeness and reproducibility, all hyperparameters employed in DoctorAgent-RL are detailed in Table 5. We observed that the decoupled clip approach from DAPO (Yu et al. 2025) significantly enhances exploration during RL training, and thus, we adopted it for our final training process.